OpenStudy (anonymous):

Pleass please help How many solutions are there to the following system of equations? 2x + 6y = 4 5x + 15y = 10 A. 0 B. 1 C. 2 D.infinitely many

OpenStudy (alexandervonhumboldt2):

2

OpenStudy (texaschic101):

5x - 15y = 10.....reduce by dividing by 5 x - 3y = 2 put in y = mx + b form -3y = -x + 2 y = 1/3x - 2/3 2x + 6y = 2 -- reduce by dividing by 2 x + 3y = 1 put it in y = mx + b form 3y = -x + 1 y = -1/3x + 1/3 since the slopes are different, then there is only 1 solution